OpenStudy (anonymous):

let f(t) be an odd function. using properties of the definite integral plus simple substitution show that if f is continuous on [-a,a] for a positive number a, then integral ∫a-af(t)dt=0

myininaya (myininaya):

\[\int\limits_{-a}^{a}f(t) dt\] we know \[f(-t)=-f(t)\] assuming a>0 we have --|---|---|--- -a 0 a btw -a and 0 we have -f(t) btw a and 0 we have f(t) \[\int\limits_{-a}^{0}f(t)dt+\int\limits_{0}^{a}f(t) dt\] \[-\int\limits_{0}^{-a}f(t)dt+\int\limits_{0}^{a}f(t) dt\] let u=-t => du=-dt \[-\int\limits_{0}^{a}f(-u) (-du)+\int\limits_{0}^{a}f(t) dt\] \[-\int\limits_{0}^{a}f(u) du+\int\limits_{0}^{a}f(t) dt\] \[-(F(a)-F(0))+(F(a)-F(0))=0\]

OpenStudy (anonymous):

omg thank you soo much